if A x B = 33541squared how can i figure out what A (length)and B (width) equal without spending an hour doing trial and error??
Answer by Edwin McCravy(20055)

Since the only whole number products that end in 1 
in the multiplication facts that you learned in 4th 
grade are

 1×1  = 1
 1×11 = 11
 3×7  = 21
 9×9  = 81
11×11 = 121

So the only numbers you need to try are numbers ending with
1,3,7,or 9

So 

you try dividing 33541 by 3.  No for you get 11180.33333
you try dividing 33541 by 7.  No for you get 4791.571429
you try dividing 33541 by 11.  No for you get 3049.181818
you try dividing 33541 by 13.  No for you get 2580.076923
you try dividing 33541 by 17.  Yes for you get 1973.

So A = 17 and B = 1973.

Edwin